Suppose demand function are                                                                                                                        

P1 = 80 – 2.5Q1

P2 = 180 – 10Q2

The Total Cost function is

TC = 50 + 40Q

find the price of two markets and amount of output to be sold in each market so that profit is maximized. Also find the total profit to be made from the strategy of price discrimination.



1
Expert's answer
2021-05-04T07:32:12-0400
"Solution"

Let's find the optimal sales to the first "q_1" and to the second "Q_2". Remember to achieve this the "MR=MC"

"TC=50+40Q\\
MC=40\\
But\ Total\ Revenue=P*Q\\
Q(80-2.5q_1)=80q-2.5q_1^2\\
Determine, MR\\
=80-5q_1\\
But\ MR=MC\\
80-5q_1=40\\
q_1=8\\
Therefore; P_1=80-2.5(8)\\
=60"


"P_2;\\
TR=P_2*Q\\
Q(180-10Q)\\
=180q-10q_2^2\\
MR=180-20q_2\\
but\ MR=MC\\
180-20q_2=40\\
q_2=7\\
Therefore;\ P_2=180-10(7)\\
=110"